OpenStudy (spring98):

Multiply (2.0 • 10^–4) • (3.1 • 10^–20). Express the answer in scientific notation. A. 6.2 • 10^–80 B. 6.2 • 10^–24 C. 6.2 • 10^24 D. 6.2 • 10^80

OpenStudy (anonymous):

It should be 6.2*10^-24

OpenStudy (spring98):

how do you know?

OpenStudy (spring98):

@tiara.c how do you know?

OpenStudy (anonymous):

(n x 10^a) • (m x 10^b) = (n · m) x 10^a+b

OpenStudy (madhu.mukherjee.946):

@Spring98 i'm sorry i could not help you

OpenStudy (madhu.mukherjee.946):

@tiara.c thanks for helping spring

OpenStudy (spring98):

it's fine i got help from @tiara.c

OpenStudy (anonymous):

and if something like this happens: (5.1 x 10^4) • (2.5 x 10^3) = 12.75 x 10^7, where the first number isn't in correct scientific notation, then you would have to move the decimal and change the exponent as well. like: 1.275 x 10^8

OpenStudy (anonymous):

You're welcome:)

OpenStudy (spring98):

Thanks @tiara.c u are very helpfull

OpenStudy (anonymous):

I'm glad I could help:)